OpenStudy (moongazer):

how do you get the domain of this:

OpenStudy (moongazer):

how do you find it? step-by-step please :)

OpenStudy (moongazer):

√(x^2 - 9x +20)/√(x^2 - 5x + 6) I already know the domain. But, how do you get it?

ganeshie8 (ganeshie8):

first workout contraints for squareroot : the thing inside squareroot cannot be negative

ganeshie8 (ganeshie8):

next work out constraints for denominator : denominator cannot be equal to 0

ganeshie8 (ganeshie8):

consider numerator first : \(\sqrt{x^2 - 9x +20}\)

ganeshie8 (ganeshie8):

can u work out domain for that ?

OpenStudy (moongazer):

{x|x>=5} is that right?

ganeshie8 (ganeshie8):

somewhat right :) lets see how to solve

ganeshie8 (ganeshie8):

x^2-9x+20 >= 0 x^2-4x-5x+20 >=0 x(x-4) -5(x-4) >= 0 (x-4)(x-5) >= 0 x <= 4 or x >= 5

ganeshie8 (ganeshie8):

thats one constraint for numerator, lets find out the constraints for denominator

OpenStudy (moongazer):

ohhhh

OpenStudy (moongazer):

I think I am starting to understand it :)

ganeshie8 (ganeshie8):

good :) since, denominator is also under radical, it must be >=0 : x^2 - 5x + 6 >= 0 x^2-3x-2x + 6 >= 0 x(x-3) -2(x-3) >= 0 (x-2)(x-3) >= 0 x <= 2 or x >= 3

ganeshie8 (ganeshie8):

with those two things, we finished with the first step of finding constraints for radicals

ganeshie8 (ganeshie8):

so far we have this : x <= 4 or x >= 5 x <= 2 or x >= 3

OpenStudy (moongazer):

yup, I understand it :)

ganeshie8 (ganeshie8):

next one is denominator can never equal to 0. so, x^2 - 5x + 6 \(\ne\) 0 x^2-3x-2x + 6 \(\ne\) 0 x(x-3) -2(x-3) \(\ne\) 0 (x-2)(x-3) \(\ne\) 0 x \(\ne\) 2 or x \(\ne\) 3

OpenStudy (moongazer):

w8 Isn't it that it should be: x >= 4 or x >= 5 x >= 2 or x >= 3 ???

ganeshie8 (ganeshie8):

lets combine all constraints and make a meaningful constraint

ganeshie8 (ganeshie8):

which one ?

OpenStudy (moongazer):

(x-4)(x-5) >= 0 (x-2)(x-3) >= 0

ganeshie8 (ganeshie8):

oh i got ur question, il give quick explanation

ganeshie8 (ganeshie8):

you comfortable with parabola graph ?

OpenStudy (moongazer):

I think so

ganeshie8 (ganeshie8):

il show u in graph why its x <=4 , x >=5

OpenStudy (moongazer):

I know few things about parabola

OpenStudy (moongazer):

ok

ganeshie8 (ganeshie8):

im sketching this parabola : (x-4)(x-5) >= 0

ganeshie8 (ganeshie8):

it intersect x axis at 4, 5 right ?

ganeshie8 (ganeshie8):